I just need help with figuring out which method to use for the below equation(s) and getting to the answer.
Solve graphically.
5x+y=-1
3x+6y=12
2x+y=-7
5x+5y=-10

Answer by ikleyn(52778)

.
I just need help with figuring out which method to use for the below equation(s) and getting to the answer.
Solve graphically.
5x+y=-1
3x+6y=12

2x+y=-7
5x+5y=-10
~~~~~~~~~~~~~~~~~~~~~~~~~

1.  They say "solve graphically" . . . .


2.  If you still want to solve algebraically then

    a) Use substitution method for the first sydtem.

    b) For the second system you may use either Subst or Elimination method.
       You may first cancel the factor 5 in both sides of the second equation.

Do not struggle !

Quickly solve everything by both methods !!


On solving systems of linear equations in two unknowns see the lessons
    - Solution of a linear system of two equations in two unknowns by the Substitution method
    - Solution of a linear system of two equations in two unknowns by the Elimination method
    - Solution of a linear system of two equations in two unknowns using determinant
    - Geometric interpretation of a linear system of two equations in two unknowns
    - Solving word problems using linear systems of two equations in two unknowns
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Happy learning !